Grade 12Electrostatics

Force of attraction between two co axial short electric dipoles

Sudharshan Jayakumar
13 Years agoGrade 12

3 Answers

Biswajit Das
11 Years ago
We know that, Coulomb force, F = Kq1q2/d2
Dipole moment, p =qd.
So, q1= p1/d and q2= p2/d
Putting these values in Coulomb force, we get,
F = K(p1/d) (p2/d)/d2
= K (p1) (p2)/d4
Therefore, Force is varies fourth power of distance.
Ajeet Tiwari
6 Years ago
hello students


We know that, Coulomb force, F = Kq1q2/d2 Dipole moment, p =qd. So, q1= p1/d and q2= p2/d Putting these values in Coulomb force, we get, F = K(p1/d) (p2/d)/d^2 = K (p1) (p2)/d^4 Therefore, Force is varies fourth power of distance.

Hope it helps
Thankyou


Yash Chourasiya
6 Years ago
Dear Students

We know that, Coulomb force, F = Kq1q2/d2
Dipole moment, p = qd.
So, q1= p1/d and q2 = p2/d
Putting these values in Coulomb force, we get,
F = K(p1/d) (p2/d)/d2
= K (p1) (p2)/d4
Therefore, It is clearly seen that Force varies with fourth power of distance.

I hope this answer will help you.
Thanks & Regards
Yash Chourasiya